OpenStudy (kaitlyn_nicole):

A Web store offers two versions of a popular song. The size of the standard version is 2.7 megabytes. The size of the high quality is 4.4 MB. Yesterday, there were 990 downloads of the song, for a total download size of 3268 MB. How many downloads of the standard version were there?

OpenStudy (kaitlyn_nicole):

@krisssssssssssss @sweetburger @Ellie201

OpenStudy (zyberg):

Try to make up a system of equations: x - standard; y - high quality; x + y = 990; 2.7x + 4.4y = 3268; Is it clear what to do next?

OpenStudy (kaitlyn_nicole):

would you plug in 0 for x and solve for y?

OpenStudy (zyberg):

@kaitlyn_nicole, no, you don't need to assume any zeroes. Just solve y in terms of x and plug that in the second equation: y = 990 - x 2.7x + 4.4(990 - x) = 3268. You should be able to solve this equation :)